Hello, I need help with a problem, does anyone know how to solve this problem !
Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
x2 = -4x – 4

Answer by stanbon(75887)

Use the discriminant to determine whether the following equations have solutions that are: two different rational solutions; two different irrational solutions; exactly one rational solution; or two different imaginary solutions.
x^2 = -4x – 4
---
x^2 + 4x + 4 = 0
---
a = 1 ; b = 4 ; c = 4
----------------------------
Discriminant: b^2-4ac = 4^2-4*1*4 = 16 - 16 = 0
----
That means there is exactly one rational solution.
=========================================================
Cheers,
Stan H.
============